OpenStudy (anonymous):

Log4 X-log16 (x+3)=1/2 The 4 and 16 are bases can someone please tell me how to solve this?

OpenStudy (amistre64):

change of base perhaps??

OpenStudy (amistre64):

4^y = x 16^y-3 = x would that work out ?

OpenStudy (anonymous):

change both bases to 4, then use the laws of exponents to combine\[\log_{16}(x+3)=y\] means \[16^y=(x+3)\] \[4^{2y}=x+3\] \[2y=log_4(x+3)\] \[y=\frac{log_4(x+3)}{2}\]

OpenStudy (amistre64):

4^y - 4^(2y-6) = 1/2

OpenStudy (amistre64):

i was close lol

OpenStudy (anonymous):

change of base gets me two joined equations with a factored out x

OpenStudy (amistre64):

log4(x) - [log4(x+3)/log4(16)] = 1/2 log4(x) - (1/2)log4(x+3) = 1/2

OpenStudy (anonymous):

idk if i can change both bases to 4

OpenStudy (amistre64):

log4(x/sqrt(x+3)) = 1/2

OpenStudy (anonymous):

you can do that?

OpenStudy (amistre64):

sqrt(4) = sqrt(x/(x+3)) i believe so :)

OpenStudy (anonymous):

then exponentiate to get \[\frac{x}{\sqrt{x+3}}=4^{\frac{1}{2}}\]

OpenStudy (amistre64):

4 = x/x+3 4x+12 = x 3x = -12 x = -4...maybe

OpenStudy (amistre64):

i got lost in it for a moment....

OpenStudy (anonymous):

no way -4 is an answer since you cannot take log of -4

OpenStudy (amistre64):

i know i know.... my brain took a momentary leave of absence lol

OpenStudy (anonymous):

the 4 is not negative

OpenStudy (amistre64):

sqrt(4) = x/(x+3) is what i meant to type :)

OpenStudy (amistre64):

4 = x^2/x^2+9+6x

OpenStudy (amistre64):

4x^2 +24x +36 = x^2 3x^2 +24x +36 = 0... thats better me thinks

OpenStudy (amistre64):

3(x^2 +8x +12) = 0

OpenStudy (amistre64):

x = -6 and -2 by that accounting...... so maybe not ;)

OpenStudy (anonymous):

oops start with \[\frac{x}{\sqrt{x+3}}=4^{\frac{1}{2}}\]

OpenStudy (amistre64):

...... yeah, it would help if I did the actual problem instead of making up me own lol

OpenStudy (anonymous):

\[x=2\sqrt{x+3}\]

OpenStudy (anonymous):

\[x^2=4(x+3)=4x+12\] \[x^2-4x-12=0\] \[(x-6)(x+2)=0\] \[x=6, x=-2\] -2 not a solution. we can try 6 to see if it works.

OpenStudy (anonymous):

yup 6 it is

OpenStudy (anonymous):

hey where did the log16 go?

OpenStudy (amistre64):

we..and be that i me sat.... modified it to a log4(x+3)/2

OpenStudy (anonymous):

I get it its just i do not know where the log16 went the log of16 is 1.20411

OpenStudy (amistre64):

change of base just changes the appearance and not the value

OpenStudy (amistre64):

log4(16) not log(16)

OpenStudy (anonymous):

let us go slow. \[log_{16}(x)=y\] means \[16^y=x\] for example \[log_{16}(4096)=3\] because \[16^3=4096\]

OpenStudy (anonymous):

but \[16=4^2\] so it is also true that \[(4^2)^3=4069\] \[4^6=4069\] so \[log_4(4069)=6\]

OpenStudy (anonymous):

so the log base 16 did not disappear, just converted to log base 4

OpenStudy (anonymous):

oho i get it thank you :]

OpenStudy (anonymous):

i got X=2.6667

OpenStudy (anonymous):

how did you get that? i am fairly certain x = 6